### v1

# Diffusiophoretic corner flows

Abstract
We study flows induced in a two-dimensional corner by the chemical activity of the confining boundaries. A catalytic reaction on the surfaces drives the diffusiophoretic flow of viscous fluid in the whole domain. We show that phoretically active sectors can generate eddies in the corner, similar to Moffatt eddies, induced mechanically in a similar confinement. We analyse the exact solution of the diffusion problem in the wedge geometry, coupled to diffusiophoretic flow generation in the slip-velocity formulation, to ultimately arrive at the stream function of the flow. We discuss cases in which simple exact solutions for the flow are available. Our work can inspire the design of microscale mixing devices involving dead-end pores in microfluidic devices and can serve as benchmarks for numerical solutions for complex geometries in (diffusio)phoretic systems.
I Introduction
The microscale manipulation of flowing fluids remains at the core of multiple modern applications, from microfluidic appliances in diagnostics and substance testing to small-scale chemical synthesis and industrial precision manufacturing technologies. In Stokes flow conditions appropriate for sub-millimetre fluidic systems, flow can be controlled globally by imposing external forces or pressure gradients that induce laminar flow with only little mixing due to molecular diffusion. While easy to control, global forcing mechanisms pose challenges for applications that require local mixing, selective pumping, or the manipulation of suspended particles within confined environments [1]. On the other hand, biological and nanorobotic actuation mechanisms, such as ciliated surfaces, rely on localised forcing to achieve swimming, pumping, mixing, nutrient capture, and sensing [2, 3]. In these settings, surface forcing becomes the key driver of macroscopic bulk flow. In confined geometry, such as in microchannels or pores, this mechanical activity is often coupled to the geometry of the flow domain, and the resulting asymmetry is responsible for creating flow.
A promising approach exploits phoretic mechanisms, in which surface-generated gradients (of concentration, temperature, etc.) induce effective slip flow on confining surfaces, that in turn gives rise to a bulk flow [4]. Diffusiophoresis, the motion of particles and fluids in response to solute concentration gradients [5, 6, 7], is now a well-established active propulsion mechanism, with numerous applications in artificial active matter [8, 9].
The presence of local chemical gradients in microfluidic channels can lead to cooperation or competition with global advective flow, enabling size-dependent colloid transport [10], or particle focussing that can be controlled by manipulating channel geometry and its chemical activity [11]. In particular, the presence of dead-end pores can be used to induce a concentration difference between the main channel and pores large enough to entrain particles [12], or capture and retain them [13], affecting filtration and dispersion in porous media [14]. Active or catalytic pores have also been proposed as local pumps and mixers capable of driving sustained fluid transport without external pressure gradients [15, 16, 17]. When considering catalytic active surfaces, geometric asymmetry alone is sufficient to create heterogeneous concentration fields that can induce propulsion [18, 19] or pumping [20, 21, 22]. Since recent advances in fabrication allow for precise spatial patterning of active regions on surfaces [23, 24], non-uniform coverage can be used together with geometric features to control microscale flow locally.
Here, we focus on a planar wedge-like geometry of dead-end pores, the walls of which are endowed with chemical activity, and which are filled with viscous fluid. Moffatt [25, 26] was the first to examine the effect of this confinement on a flow that emerges in response to a disturbance that acts far away from the tip of the wedge, as well as from a mechanically active sector on the surface, imposing a slip flow on the boundary. In both cases, the celebrated Moffatt eddies emerge as a solution, with an infinite sequence of vortices being created in the fluid, as later seen experimentally by Taneda [27]. Self-similar vortical solutions emerge frequently in externally forced confined flows, such as wedge-shaped trenches with a free surface [28], cone-like geometry [29], electrohydrodynamic flows [30], simulations of driven cavity flows [31, 32], and in ice flows over subglacial mountain valleys [33]. The classical problem of flow actuation by moving boundaries is perhaps best illustrated in the context of corner flows by the Taylor’s scraper problem [34], where one boundary moving boundary ’scrapes’ the fluid, causing its outward motion along the immobile wall. Moffatt’s analysis of corner flow driven by a partial slip on the walls [26], akin to conveyor belts shearing the fluid locally, is an example of (local) mechanical actuation. Here, we explore a similar concept with a chemical actuation mechanism that couples to flow through diffusiophoresis.
In this work, we consider a non-uniform coverage of the wedge with a catalyst that induces the release or capture of solute. The heterogeneous concentration field, emerging from geometric asymmetry of the fluid domain, drives slip flow on the active surfaces, which in turn induces bulk flow that takes the form of a sequence of vortices. We calculate the flow analytically in the simplest cases. Next, using the Mellin transform formalism, we present an approach that applies to any coverage of the walls with chemical activity.
The paper is structured as follows. First, in Sec. II, we present the general mathematical framework of diffusiophoretic Stokes flows, describe the geometry of the problem, and the relevant physical quantities. In Sec. III, we discuss the general solutions of the diffusion equation for the solute and the biharmonic equation for the flow stream function in the wedge geometry. We also introduce the formalism of Mellin transforms suited to the geometry considered. Next, in Sec. IV we present analytical solutions for the diffusiophoretic flows induced by uniform coverage of one or both walls with a catalyst. In the following Sec. V we discuss the solute concentration field emerging with a single active sector on one of the walls and the other boundary being absorbing, for different geometric settings. The complementary problem of diffusion with an active sector and a reflective boundary is treated in Sec. VI. Finally, we show in Sec. VII how the solute concentration fields translate to the flow, and obtain the flow field numerically by solving for the stream function. Additionally, in Sec. VIII we discuss the case of the chemical activity of the walls given by analytical functions rather than patches, represented by step-like profiles. We conclude the paper in Sec. IX.
II Phoretic flow generation
II.1 General framework
We adopt a continuum description of phoretic transport, following established theoretical frameworks  [5, 6, 7], to analyse the two-dimensional flow field generated within a wedge formed by two semi-infinite lines starting from one point. The fluid within the wedge is characterised by dynamic viscosity $\eta$, density $\rho$, and contains solute of local concentration $C$, with diffusivity $\kappa$.
The chemical activity $\mathcal{A}$ of a surface $\mathcal{S}$ quantifies the fixed rate of solute release ($\mathcal{A}>0$) or absorption ($\mathcal{A}<0$) on the surface by
$\kappa\bm{n}\cdot\nabla C=-\mathcal{A}\quad\textrm{ on }\mathcal{S}.$ (1)
Due to the short-range interaction of solute molecules with the cavity boundary, local concentration gradients result in the motion of the solute and, consequently, drive the motion of the fluid [4]. Assuming that the thickness of the interaction layer is small compared to the dimensions of the cavity, the classical slip-velocity formulation can be used [35], and local solute gradients induce an effective slip velocity on the boundaries
$\bm{u}=\mathcal{M}(\bm{1}-\bm{n}\bm{n})\cdot\nabla C\quad\textrm{ on }\mathcal{S},$ (2)
which drives the bulk motion of the fluid. Here, $\mathcal{M}$ is the local phoretic mobility at the surface $\mathcal{S}$. It is related to the surface-solute interaction potential [4].
Introducing $\mathcal{R}$ as the characteristic length scale, and $\mathcal{U}=|\mathcal{AM}|/\kappa$ as the characteristic phoretic velocity generated along $\mathcal{S}$, we can define the Péclet and Reynolds numbers
$\mathrm{Pe}=\frac{\mathcal{UR}}{\kappa},\qquad\text{Re}=\frac{\rho\,\mathcal{U}\mathcal{R}}{\eta}.$ (3)
which quantify, respectively, the relative importance of solute advection and diffusion as transport mechanisms and effects of inertia and viscosity in forces that shape the flow. When Pe is small enough, the solute dynamics are purely diffusive and are governed by Laplace’s equation
$\nabla^{2}C=0,$ (4)
in the whole domain. Provided that inertial effects are negligible (i.e. $\text{Re}\ll 1$), the flow and pressure fields satisfy the incompressible Stokes equations:
$\eta\nabla^{2}\bm{u}$ $=\nabla p,$ (5)
$\nabla\cdot\bm{u}$ $=0.$ (6)
The diffusive Laplace’s problem for the solute concentration $C$ effectively decouples from the hydrodynamic problem and may be solved independently. The concentration $C$ can be used to compute the slip flow along $\mathcal{S}$, eq. (2), which is the boundary condition for the flow field in eq. (5) within the cavity.
Given the boundary conditions that we assume in this problem, and the governing Laplace’s equation, if $C_{0}$ is a constant and the concentration field $C$ is a solution, then $c=C+C_{0}$ is a solution too. The boundary conditions (1), (2) do not depend on the absolute value of the local concentration. For all $C_{0}$, the resulting flow will be the same. This also implies that $c$ can be negative. After shifting by a positive $C_{0}$, we will still obtain a physical solution. In the next sections, we will therefore use the relative concentration $c$.
II.2 Nondimensionalization of the model
Having defined $\mathcal{U}$ and $\mathcal{R}$ as the characteristic velocity and length, respectively, the characteristic concentration is $|\mathcal{A}|\mathcal{R}/\kappa$ and the characteristic pressure is $\eta|\mathcal{AM}|/\mathcal{R}\kappa$. The dimensionless activity and mobility are given by $A=\mathcal{A}/|\mathcal{A}|$ and $M=\mathcal{M}/|\mathcal{M}|$.
The nondimensional governing equations are as follows. For the diffusion of solute, we obtain
$\nabla^{2}c$ $=0,$ (7a)
$\bm{n}\cdot\nabla c$ $=-A\quad\textrm{ on }\mathcal{S},$ (7b)
while the flow problem becomes
$\nabla^{2}\bm{u}$ $=\nabla p,$ (7c)
$\nabla\cdot\bm{u}$ $=0,$ (7d)
$\bm{u}$ $=M(\bm{1}-\bm{n}\bm{n})\cdot\nabla c\quad\textrm{ on }\mathcal{S}.$ (7e)
In the following, we apply this general framework to the specific geometry of a narrowing corner to explore the ways in which fluid motion can be actuated within such a cavity by purely chemical means, with no moving mechanical parts.
III Corner flows
The problem of phoretic flow generation in wedge qualitatively resembles that leading to Moffatt eddies [25, 26]. In the classical problem analysed by Moffatt [25], an infinite sequence of eddies is formed in a wedge-shaped planar domain filled with viscous fluid due to a disturbance acting at asymptotically large distances from the corner. Eddies emerge as a self-similar solution to a second-kind eigenvalue problem for the streamfunction. However, similar flow structures can arise when the fluid is actuated by a moving boundary, as in the famous Taylor’s scraper problem [34], where one of the walls of the wedge-shaped cavity moves with constant speed and drives macroscopic flow. A refined variant of this geometric setting involves motion of a portion of the boundary, when a section of it is endowed with slip velocity, as explored by Moffatt [26]. Using Mellin transform techniques, Moffatt found a solution of this problem for a single moving region (and two regions placed symmetrically), akin to transmission belts mounted within the walls. The solution again involves a sequence of corner eddies. This mechanical example operates using the same principle as one implemented in our case.
Here, we focus on a similar geometry but allow the walls to exhibit chemical activity, which drives the flow. The diffusiophoretic flow generation mechanism leads, in general, to a non-uniform slip velocity profile on the bounding walls. However, for the case of uniform coverage of either one or both walls with catalyst, the induced slip velocity remains constant, therefore directly reducing to the previously obtained results. As we demonstrate analytically in the following, both antisymmetric and symmetric flows can be induced, depending on the specific coverage pattern.
III.1 Solute concentration
We first focus on the problem of diffusion of a solute in the corner domain. Upon introducing cylindrical polar coordinates $(\rho,\theta)$ (see figure 1), the stationary Laplace’s equation for the concentration field $c$ takes the form
${\rho}\frac{\partial}{\partial\rho}\left(\rho\frac{\partial c}{\partial\rho}\right)+\frac{\partial^{2}c}{\partial\theta^{2}}=0.$ (8)
Suppose that the wedge extends in polar coordinates from $\theta=0$ to $\theta=\alpha$. The boundary conditions are imposed on the normal gradient of concentration, $\bm{n}\cdot\nabla c$, and are thus given by the chemical activity distribution on the walls, $A_{0}(\rho)$ and $A_{\alpha}(\rho)$ as
$\frac{1}{\rho}\frac{\partial c}{\partial\theta}=-A_{0}(\rho),$ (9)
if the condition is imposed on the wall $\theta=0$, and
$\frac{1}{\rho}\frac{\partial c}{\partial\theta}=A_{\alpha}(\rho),$ (10)
if the condition is imposed on the wall $\theta=\alpha$.
Figure 1: Geometry of the diffusiophoretic corner flow setup. An active patch, marked on the $\theta=0$ wall, covering the radial section $\rho\in[a,b]$, releases solute (grey arrows) and generates an inhomogeneous concentration field which drives circulatory flow indicated by schematic streamlines.
III.2 Stokes flow
Once the concentration distribution $c(\rho,\theta)$ is known, the slip flow velocity $\bm{u}_{s}=u_{s}\bm{e}_{\rho}$ on the boundary is determined as
$u_{s}=M\left.\frac{\partial c}{\partial\rho}\right|_{\mathcal{S}},$ (11)
which becomes the boundary condition for a two-dimensional flow in the fluid domain. In this problem, it is convenient to introduce the Stokes stream function $\Psi(\rho,\theta)$ that obeys a biharmonic equation
$\nabla^{4}\Psi=0,$ (12)
with the boundary conditions on the two planes $\theta=\{0,\alpha\}$ being
$u_{\rho}=\frac{1}{\rho}\frac{\partial\Psi}{\partial\theta}=u_{s},\qquad u_{\theta}=-\frac{\partial\Psi}{\partial\rho}=0,\qquad\text{on {\hbox{\theta= \{0 , \alpha\}}}.}$ (13)
III.3 Solution of Laplace and biharmonic equations in the Mellin space
We shall now introduce the formalism of Mellin transforms to solve the more general case of a single active sector. Indeed, Melling transforms have become the traditional method for solving boundary-value problems in wedge-shaped regions [36, 25, 37]. In cylindrical coordinates, Laplace’s equation takes the form of Eq. (8). Let $\mathds{1}_{\Omega}$ be an indicator function of a set $\Omega$. The Mellin transform $\mathscr{M}_{p}$ of concentration $c$ is denoted as $\bar{c}$
$\mathscr{M}_{p}\{c(\rho,\theta)\}:=\int_{0}^{\infty}\rho^{p-1}c(\rho,\theta)\,d\rho=:\bar{c}(p,\theta).$ (14)
In addition, $\mathscr{M}^{-1}$ is the inverse Mellin transform, which reads
$c(\rho,\theta)=\mathscr{M}^{-1}\{\bar{c}(p,\theta)\}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}\bar{c}(p,\theta)\rho^{-p}dp.$ (15)
The condition for the existence of an inverse transform is for $\int_{0}^{\infty}\rho^{\gamma-1}c(\rho,\theta)\,d\rho$ to exist. Unless stated otherwise, $\gamma=0$. Then, for appropriate $c(\rho,\theta)$,
$\mathscr{M}_{p}\{\rho\partial_{\rho}c(\rho,\theta)\}=\mathscr{M}_{p+1}\{\partial_{\rho}c(\rho,\theta)\}=-p\bar{c}(p,\theta),$ (16)
where we introduced an abbreviation for derivatives, $\partial_{\rho}\equiv\partial/\partial\rho$, etc. The Laplace’s equation for concentration, (8), implies that $p^{2}\bar{c}+\partial_{\theta}\bar{c}=0$ and for some $C(p)$, $D(p)$ we have
$\bar{c}(p,\theta)=C(p)e^{ip\theta}+D(p)e^{-ip\theta}.$ (17)
We now turn to the flow problem. Let $\Psi$ be a stream function. The biharmonic equation, $\nabla^{2}\nabla^{2}\Psi=0$, in cylindrical coordinates and after applying the Mellin transform $\mathscr{M}_{p+4}$ becomes
$\left\{\partial_{\theta}^{4}+[(p+2)^{2}+p^{2}]\partial_{\theta}^{2}+p^{2}(p+2)^{2}\right\}\bar{\Psi}(p,\theta)=0,$ (18)
where we write $\bar{\Psi}(p,\theta):=\mathscr{M}_{p}\{\Psi(\rho,\theta)\}$. The general solution then reads
$\bar{\Psi}=F_{1}(p)\cos(p\theta)+F_{2}(p)\sin(p\theta)+G_{1}(p)\cos((p+2)\theta)+G_{2}(p)\sin((p+2)\theta).$ (19)
IV Phoretic walls with uniform coverage
To study the effect of chemical activity on corner flows, we first focus on the case when the coverage by catalyst is uniform, and the heterogeneity of the concentration field stems purely from the geometry of the wedge. In the two examples studied in the following, either one or both walls are active, and the problem admits analytical solutions.
IV.1 A single active wall
In the first simple example, we consider a single active wall, with constant activity $A$ on the surface. To ensure the existence of a steady-state solution, we assume that the other boundary is absorbing, and thus we can write the boundary conditions for the concentration problem as
$\frac{\partial c}{\partial\theta}=-A\rho\qquad$ $\text{for $\theta=0$},$ (20)
$c=0\qquad$ $\text{for $\theta=\alpha$}.$ (21)
Using the separated form $c=R(\rho)\Theta(\theta)$ as an ansatz, we find the general solution
$R(\rho)=a\rho^{\lambda}+b\rho^{-\lambda},\qquad\Theta(\theta)=p\cos\lambda\theta+q\sin\lambda\theta.$ (22)
Applying the boundary conditions, we obtain the concentration field as
$c(\rho,\theta)=-\frac{A\rho\sin(\theta-\alpha)}{\cos\alpha},$ (23)
from which the tangential gradient dictates the slip velocity distribution on the surface as
$\bm{u}=MA\tan\alpha\ \bm{e}_{\rho}\qquad$ $\text{for $\theta=0$},$ (24)
$\bm{u}=\bm{0}\qquad$ $\text{for $\theta=\alpha$}.$ (25)
With this velocity profile, the problem becomes identical to the famous Taylor’s scraper [34], with the imposed slip velocity $U=MA\tan\alpha$. The stream function then has the form
$\Psi=U\rho f(\theta),$ (26)
where $f(\theta)$ satisfies an ordinary differential equation
$f{{}^{\prime\prime\prime\prime}}+2f{{}^{\prime\prime}}+f=0,$ (27)
with the boundary conditions $f(0)=0$, $f^{\prime}(0)=1$, $f(\alpha)=0$, $f^{\prime}(\alpha)=0$. This is a particular form of the general solution to planar elasticity problems for the biharmonic Airy stress function due to [38]. The resulting stream function reads:
$\Psi=U\rho f(\theta)=\frac{U\rho}{\alpha^{2}-\sin^{2}\alpha}\left[\alpha(\alpha-\theta)\sin\theta-\theta\sin\alpha\sin(\alpha-\theta)\right].$ (28)
The resulting flow fields are drawn in figure 2 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is asymmetric, reflecting the boundary conditions.
Figure 2: Phoretic flow induced by the activity of one phoretic wall in a wedge of angle $\alpha$ for (left to right) $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. Flow streamlines are marked in white. The colour map indicates the total velocity magnitude.
IV.2 Two phoretic walls
A simple generalisation of the problem of one wall involves covering both walls of the wedge with catalyst, thereby imposing a slip velocity symmetrically for both $\theta=0$ and $\theta=\alpha$. In this case, we expect the flow to be symmetrical about the wedge bisector.
The concentration profile satisfying the fixed-flux boundary conditions at both walls, namely,
$\frac{\partial c}{\partial\theta}=-A\rho\quad$ $\textrm{for}\quad\theta=0$ (29a)
$\frac{\partial c}{\partial\theta}=A\rho\quad$ $\textrm{for}\quad\theta=\alpha,$ (29b)
follows directly as
$c(\rho,\theta)=A\rho\frac{\sin(\theta-\alpha)-\sin\theta}{1-\cos\alpha}.$ (30)
Equation (27) remains valid with the new boundary conditions $f(0)=f(\alpha)=0$ and $f^{\prime}(0)=f^{\prime}(\alpha)=1$. The stream function in this case takes on the form
$\Psi=\frac{U\rho}{\alpha-\sin\alpha}\left[(\alpha-\theta)\sin\theta-\theta\sin(\alpha-\theta)\right].$ (31)
Figure 3: Phoretic flow induced by the activity of two phoretic walls in a wedge of angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$.
The resulting flow fields are drawn in figure 3 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is symmetrical, with the fluid being dragged in from the sides and pushed away along the bisector direction.
V The diffusion problem with an active and absorbing wall
To explore the flow in a more general case, we focus on a corner with a single active sector of length $\ell=b-a$ placed on one of the boundaries at a distance $a$ (between $\rho=a$ and $\rho=b$), as depicted in figure 1. The problem then, contrary to the previous cases, has two natural length scales: $\ell$ and $a$.
The concentration field satisfies the Laplace’s equation, $\nabla^{2}c=0$, with the boundary conditions
$\frac{\partial c}{\partial\theta}=-A\rho\qquad$ $\text{for $\theta=0$ and $a<\rho<b$},$ (32)
$c=0\qquad$ $\text{on other boundaries}.$ (33)
V.1 Solution of the diffusion problem for a single catalytic segment
We consider a single catalytic segment $[a,b]$ placed on one of the walls ($\theta=0$), as sketched in figure 1, and analyse the solution in several specific cases.
V.1.1 The case of $a\neq 0$ and $b<\infty$
Let the wall with $\theta=\alpha$ be ideally absorptive. Then, the boundary conditions take the form
$\frac{\partial c}{\partial\theta}$ $=-A\rho\mathds{1}_{[a,b]}$ $\textrm{for}\quad\theta=0$ (34a)
$c$ $=0$ $\textrm{for}\quad\theta=\alpha.$ (34b)
It is assumed that $0<\alpha<2\pi$. The solution in Mellin space is
$\bar{c}(p,\theta)=f_{b}-f_{a},\quad\textrm{where}\quad f_{a}(p,\theta)=A\frac{a^{p+1}}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}.$ (35)
It has singularities at $p^{\prime}_{k}$ such that $p^{\prime}_{k}\alpha=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, and also at $p_{0}=-1$, and a removable singularity at $p=0$.
The concentration is then written as $c=I_{b}-I_{a}$, where
$I_{a}=\frac{1}{2\pi i}\int_{-i\infty}^{i\infty}A\frac{a}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}\left(\frac{a}{\rho}\right)^{p}dp.$ (36)
For $a\leq\rho$, the integral can be evaluated by taking a square contour over the half-plane $\text{Re}(p)>0$; for $a>\rho$, the contour is taken over $\text{Re}(p)<0$. The choice ensures that the integral over three sides of the square contour vanishes.
The poles are listed as
$p_{k}=\begin{cases}{\alpha^{-1}\left(\frac{\pi}{2}+(k-1)\pi\right)}&\text{for $k\in\mathbb{Z}_{+}$}\\ -1&\text{for $k=0$}\\ \alpha^{-1}{\left(-\frac{\pi}{2}+(k+1)\pi\right)}&\text{for $k\in\mathbb{Z}_{-}$}.\end{cases}$ (37)
Under the assumption that, for all $k$, $({\frac{\pi}{2}+k\pi})/{\alpha}\neq-1$ (i.e. $\alpha\neq{\pi}/{2},\pi,{3\pi}/{2}$) the residues are
for $k\neq 0:\quad\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}=-\frac{Aa}{\alpha p_{k}(p_{k}+1)}\cos(p_{k}\theta)\left(\frac{a}{\rho}\right)^{p_{k}},$ (38a)
for $k=0:\quad\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}=A\frac{\sin(\alpha-\theta)}{\cos\alpha}\rho.$ (38b)
Depending on the ratio of ${a}/{\rho}$, different residues are contained in the integration contour, and the orientation of the contour is different, hence
$I_{a}(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$}\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise,}\end{cases}$ (39)
with the full solution being
$c(\rho,\theta)=I_{b}(\rho,\theta)-I_{a}(\rho,\theta).$ (40)
Note that the residue at p=1p=-1 is responsible for satisfying the condition (34a). In the special cases of $\alpha=\frac{\pi}{2},\pi,\frac{3\pi}{2}$ solutions can be obtained by taking, for example, the limit of $\alpha\to\frac{\pi}{2}$. The solutions for $\alpha=\frac{\pi}{2}-\epsilon$ and $\alpha=\frac{\pi}{2}+\epsilon$ can then be compared to check if the desired accuracy has been achieved for a chosen $\epsilon$.
V.1.2 The case of $a=0$, $b<\infty$ i.e. sector $[0,b]$
The solution is obtained similarly to the previous case as
$c(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{for $b>\rho$},\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{otherwise.}\end{cases}$ (41)
V.1.3 The case of $a\neq 0$ and $b=\infty$ i.e. sector $a,\infty[$
In this case existence of Mellin transform requires that $\text{Re}(p)\leq-1$.
We thus take $\gamma=-1-\epsilon>-{\pi}/{2}$ and write the solution as
$c(\rho,\theta)=\begin{cases}-\sum_{k\leq-1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$},\\ \sum_{k\geq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise.}\end{cases}$ (42)
A simple computation shows that, in all the cases above, $c(\rho,\theta)$ with an appropriate choice of $\gamma$, satisfies the condition for the existence of the inverse transform.
In figure 4, we sketch the solutions for the concentration field for a chosen wedge opening angle $\theta=\pi/6$ in three representative cases. The catalytic sector is marked in red, and the colours indicate the absolute value of solute concentration. In all cases, the resulting concentration field is heterogeneous, which offers the possibility to drive slip flow along the active boundary.
Figure 4: Solute concentration fields for an ideally absorptive wall at $\theta=\pi/6$, a phoretic wall at $\theta=0$ with a catalytic sector at $(a,b)=\{(1,3),(0,3),(3,\infty)\}$ marked in red. Here, we assumed $A=1$ and the plotted radius of the wedge is $\rho<4$. The scale bar for the absolute concentration field, $|c|$, is common for plots (a), (b) and different for panel (c).
VI The diffusion problem with an active and a reflective wall
VI.1 Solution for a single catalytic sector with reflective walls
Consider a setting in which the wall at $\theta=0$ contains a catalytic sector with activity $A$ and the wall at $\theta=\alpha$ is not active, thus it is a reflective wall, with no-flux boundary condition. We denote the corresponding solute concentration field by $c_{0}$. The boundary conditions are then
$\frac{\partial c_{0}}{\partial\theta}=-A\rho\mathds{1}_{[a_{1},b_{1}]}\quad\textrm{for}\quad\theta=0$ (43a)
$\frac{\partial c_{0}}{\partial\theta}=0\quad\textrm{for}\quad\theta=\alpha.$ (43b)
We again assume that $0<\alpha<2\pi$. A solution satisfying these boundary conditions in the Mellin space reads
$\bar{c}_{0}(p,\theta)=-A\frac{b_{1}^{p+1}-a_{1}^{p+1}}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}.$ (44)
If $c_{\alpha}$ is a solution for the active sector with activity $A$ on the wall $\theta=\alpha$ i.e., for boundary conditions
$\frac{\partial c_{\alpha}}{\partial\theta}=0\quad\textrm{for}\quad\theta=0$ (45a)
$\frac{\partial c_{\alpha}}{\partial\theta}=A\rho\mathds{1}_{[a_{1},b_{1}]}\quad\textrm{for}\quad\theta=\alpha,$ (45b)
then the corresponding solution takes the form
$\quad\bar{c}_{\alpha}(p,\theta)=\bar{c}_{0}(p,\alpha-\theta)=-A\frac{b_{1}^{p+1}-a_{1}^{p+1}}{p(p+1)}\frac{\cos(p\theta)}{\sin(p\alpha)}.$ (46)
Let us introduce the notation
$g_{0,a}(p,\theta)$ $=-A\frac{a}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}\left(\frac{a}{\rho}\right)^{p},$ (47)
$I_{0,a}$ $=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}g_{0,a}dp.$ (48)
Applying the inverse Mellin transform to the solution (44), we can write
$c_{0}=I_{0,b_{1}}-I_{0,a_{1}}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}(g_{0,b_{1}}-g_{0,a_{1}})dp.$ (49)
We now note that $g_{0,a}$ has poles at $p^{\prime}_{k}={k\pi}/{\alpha}$, with $k\in\mathbb{Z}$, and at p=1p=-1. Under the assumption that for all $k$: $p^{\prime}_{k}\neq-1$ (that is, $\alpha\neq\pi$) the residues are
for $k\neq 0:\quad\mathop{\mathrm{Res}}_{p=p_{k}}g_{a}=-\frac{Aa}{\alpha}\frac{\cos(p_{k}(\alpha-\theta))}{p_{k}(p_{k}+1)}\left(\frac{a}{\rho}\right)^{p_{k}},$ (50a)
for $k=0:\quad\mathop{\mathrm{Res}}_{p=p_{k}}g_{a}=-\frac{Aa}{\alpha}\left[\log\left(\frac{b}{\rho}\right)-1\right],$ (50b)
for $p=-1:\quad\mathop{\mathrm{Res}}_{p=-1}g_{a}=-A\rho\frac{\cos(\alpha-\theta)}{\sin\alpha}.$ (50c)
The choice of $\gamma$ such that $\textrm{min}\{p^{\prime}_{k}:p^{\prime}_{k}>0\}=\frac{\pi}{\alpha}>\gamma>0$ guarantees that the Mellin transform of $c_{0}$ exists. Integrating over a square contour in one of the half-planes, we obtain
$I_{0,a}(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}&\text{for $a>\rho$},\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}-\mathrm{Res}_{p=-1}g_{a}&\text{otherwise.}\end{cases}$ (51)
Using Eq. (49), $c_{0}$ can now be computed straightforwardly. The total activity of the walls not being zero does not cause a contradiction, since there are no constraints on activity at $\rho=\infty$ and the solute can be absorbed or emitted at infinity. The case $\alpha=\pi$ can be solved in Cartesian coordinates or by taking the limit of $\alpha\to\pi$ in the above solution.
VI.2 Diffusion for multiple catalytic sectors
By linearity, the solution can be obtained by superposition of solutions for single sectors. If sectors $a_{i},b_{i}]$ have activity $A_{i}$, the condition of equilibrium is
$\sum_{i}A_{i}(b_{i}-a_{i})=0,$ (52)
where the sum is taken over all active sectors. This is also the condition for the solution to be physical (solute conservation). That is, unless we allow the solute to be released/absorbed at $\rho=\infty$. It can be seen that the resulting solutions are indeed continuous. Note that the sectors can overlap, so step activity profiles can be created to mimic arbitrary activity functions $A_{0}(\rho)$ and $A_{\alpha}(\rho)$ of the walls.
An example of a setting involving multiple sectors is presented in figure 5 for two active patches of opposite activity on each of the walls, and for different wedge opening angles. The activity of patches closer to the tip of the wedge is diminished by their mutual influence, and the concentration fields produced by patches further away from the confinement are more pronounced. This demonstrates the intricate interaction between the geometry of the wedge and the chemical activity of its walls.
Figure 5: Solute concentration fields for wedges with multiple active sectors. (a)-(d) The wedge angles are $\alpha=\{4\pi/7,3\pi/7,5\pi/6,\pi/6\}$. There are two active sectors of opposite activity $|A|=1$ on each wall. They span the sections $\rho\in(0.5,1.5)$ and $\rho\in(2.5,3.5)$. The plotted radius of the wedge is $\rho<4$. Patches of positive activity are marked in red, while those of negative activity are marked in blue. Panels (a)-(c) share the scale bar placed in the middle of the figure, while the scale bar for panel (d) is in the top right corner.
VI.3 The diffusion problem for an arbitrary catalytic wall
Let us now consider a setting where the wall at $\theta=0$ has arbitrary activity $A=A(\rho)$ and the wall at $\theta=\alpha$ is reflective. We will approach this problem using the Green’s function, because the Stokes equation is linear. Let $c^{G}_{0}(\rho,\rho^{\prime},\theta)$ be the Green’s function for activity $\delta(\rho-\rho^{\prime})$. Boundary conditions take the following form,
$\frac{\partial c^{G}_{0}}{\partial\theta}(\rho,\rho^{\prime},\theta)=-\rho\delta(\rho-\rho^{\prime})\quad\textrm{for}\quad\theta=0$ (53a)
$\frac{\partial c^{G}_{0}}{\partial\theta}(\rho,\rho^{\prime},\theta)=0\quad\textrm{for}\quad\theta=\alpha.$ (53b)
The solution in the Mellin space reads
$\bar{c}^{G}_{0}(p,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\rho^{\prime p},$ (54)
and the general solution for the transformed concentration becomes
$\bar{c}_{0}(p,\theta)=\int_{0}^{\infty}A(\rho^{\prime})\bar{c}^{G}_{0}(p,\rho^{\prime},\theta)d\rho^{\prime}.$ (55)
Residues of $\bar{c}^{G}_{0}$ are $p_{k}={k\pi}/{\alpha}$. We now change the order of integration in the back-transformed concentration field to arrive at
$c_{0}(\rho,\theta)=\frac{1}{2\pi i}\int_{0}^{\infty}d\rho^{\prime}A(\rho^{\prime})\int_{\gamma-i\infty}^{\gamma+i\infty}-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\left(\frac{\rho^{\prime}}{\rho}\right)^{p}dp,$ (56)
where $\gamma$ has to satisfy the same condition as in the previous section. Defining
$f(p,\rho,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\left(\frac{\rho^{\prime}}{\rho}\right)^{p},$ (57)
we evaluate the following residues
$\mathrm{Res}_{p=p_{k}}f=\begin{cases}-\frac{1}{\alpha}\frac{\cos(p_{k}(\alpha-\theta))}{p_{k}}\left(\frac{\rho^{\prime}}{\rho}\right)^{p}&\text{for $k\neq 0$},\\ -\frac{1}{\alpha}\log\left(\frac{\rho^{\prime}}{\rho}\right)&\text{for $k=0$}.\end{cases}$ (58)
Choosing the contours of integration the same way as in the previous section, we finally obtain
$c_{0}(\rho,\theta)=\int_{0}^{\rho}d\rho^{\prime}A(\rho^{\prime})\sum_{k\leq 0}\mathrm{Res}_{p_{k}}f(p,\rho,\rho^{\prime},\theta)-\int_{\rho}^{\infty}d\rho^{\prime}A(\rho^{\prime})\sum_{k>0}\mathrm{Res}_{p_{k}}f(p,\rho,\rho^{\prime},\theta).$ (59)
We note that for two arbitrary catalytic walls, superposition of solutions can be used. The condition to satisfy solute conservation is
$\int_{0}^{\infty}(A_{0}(\rho)+A_{\alpha}(\rho))d\rho=0,$ (60)
where $A_{0}(\rho)$ and $A_{\alpha}(\rho)$ are the activities of the walls $\theta=0$ and $\theta=\alpha$, respectively. It is worth noting that, for practical purposes, approximating the solution, for a given function $A(\rho)$, by making a superposition of solutions for constant activity sectors could potentially be more efficient than truncating the above series and evaluating the integrals.
VII Solution of the biharmonic equation for given solute concentration field
Having resolved the concentration field in various cases and for different boundary conditions on each wall, we now turn to the calculation of the corresponding flow field. The boundary conditions for the stream function $\Psi(\rho,\theta)$, that satisfies the biharmonic equation (12), are
$u_{\theta}=-\frac{\partial\Psi}{\partial\rho}=0\quad\textrm{for}\quad\theta=0,\alpha$ (61a)
$u_{\rho}=\frac{1}{\rho}\frac{\partial\Psi}{\partial\theta}=M\frac{\partial c}{\partial\rho}\quad\textrm{for}\quad\theta=0,\alpha,$ (61b)
indicating that the only source of flow is the slip flow at the boundary, dictated by the heterogeneous solute concentration profile. Applying the Mellin transform $\mathscr{M}_{p}$ to these conditions yields
$\bar{\Psi}=0\quad\textrm{for}\quad\theta=0,\alpha$ (62a)
$\frac{\partial\bar{\Psi}}{\partial\theta}=-pM\bar{c}\quad\textrm{for}\quad\theta=0,\alpha.$ (62b)
Using the general form of the solution for $\bar{\Psi}$, given by Eq. (19), we may now identify the coefficients to be
$G_{1}$ $=-F_{1}\hskip 210.0pt$ (63a)
$\begin{bmatrix}F_{1}\\ F_{2}\\ G_{2}\end{bmatrix}$ $=\frac{-pM}{2D(p)}\bar{c}(p,0)\begin{bmatrix}(p+1)\sin(2\alpha)-\sin(2\alpha(p+1))\\ (p+1)\cos(2a)+\cos(2a(p+1))-p-2\\ -p+(p+1)\cos(2\alpha)-\cos(2\alpha(p+1))\end{bmatrix}$+
$\phantom{=}+\frac{-pM}{2D(p)}\bar{c}(p,\alpha)\begin{bmatrix}(p+2)\sin(\alpha p)-p\sin(\alpha(p+2))\\ -((p+2)(\cos(\alpha p)-\cos(\alpha(p+2))))\\ p(\cos(\alpha p)-\cos(\alpha(p+2)))\end{bmatrix}$, (63b)
where $D(p)=(p+1)^{2}\cos(2\alpha)-p(p+2)-\cos(2\alpha(p+1))$. The way to obtain $\bar{c}(p,0)$ and $\bar{c}(p,\alpha)$ for an arbitrary setting of catalytic sectors on both walls has been demonstrated in previous sections. The remaining task is thus to invert the Mellin transform; this turns out to be challenging analytically, since zeros of $D(p)$ are difficult to obtain, unlike when inverting the Mellin transform for the concentration. We thus compute the integral (see Eq. (15)) numerically. In an illustrative case, where the active sectors on the walls span from the corner to $\rho=1$, the calculated streamlines (the isolines of $\Psi(\rho,\theta)$) can be seen in figure 6. A counterclockwise eddy is created in the immediate vicinity of the active patches, and its rotation drives a larger eddy in the opposite direction further away from the corner. We note that this flow profile is obtained using only two active patches. When combined into more complicated coverage patterns, the flow can be tuned to a particular application, e.g. localised mixing.
Figure 6: (a) Isolines of the stream function $\psi$ of diffusiophoretic corner flow for $\theta=\pi/3$, with active sectors at $\rho\in(0,1)$, emitting at $\theta=0$ and absorbing at $\theta=\alpha$. The flow in the corner eddy (yellow) is counter-clockwise and drives clockwise rotation of another eddy further away from the corner. Colours code the magnitude of $\psi$. (b) Transversal velocity profile $v_{\theta}(\rho)$, on the bisector angle of the wedge. Roots of the velocity indicate the centres of vortices. The slight wiggle for small $\rho$ is likely a numerical artifact.
VIII The diffusion problem with analytic Neumann boundary conditions
Here, we present an auxiliary result for the diffusion problem when the activity patterns are given by analytic functions, rather than step-like profiles discussed earlier. Let $A^{1}(\rho)$ and $A^{2}(\rho)$ represent the (analytic) distribution of activity on the walls with $\theta=0$ and $\theta=\alpha$, respectively. We make the following ansatz for the concentration profile $c$ in the fluid domain
$c=\lambda+\sum_{n\geq 1}(F_{n}\cos(n\theta)+G_{n}\sin(n\theta))\rho^{n},$ (64)
and expand $A^{1}$ and $A^{2}$ in power series in $\rho$ as
$A^{1}=\sum_{n\geq 1}A^{1}_{n}\rho^{n},\ A^{2}=\sum_{n\geq 1}A^{2}_{n}\rho^{n}.$ (65a)
Imposing the fixed flux boundary conditions, Eqs. (9) and (10), we arrive at linear equations for the coefficients of the expansion in Eq. (64). Solving these equations, we obtain
$F_{n}$ $=-\frac{A^{1}_{n-1}\cos(n\alpha)+A^{2}_{n-1}}{n\sin(n\alpha)},$ (66a)
$G_{n}$ $=-\frac{A^{1}_{n-1}}{n},$ (66b)
for $n\geq 1$. While $\lambda$ remains arbitrary, it can be fixed by an additional boundary condition, such as the value of $c(\rho=0)$. Then, we get $c(\rho=0)=\lambda$. The condition for existence of $F_{n}$ and $G_{n}$ such that boundary conditions can be satisfied by this ansatz is $\sin(n\alpha)\neq 0$ (i.e. $\alpha\neq q\pi$, where $q\in\mathbb{Q}^{+}$). If $\sin(n\alpha)=0$, a more general ansatz is required. Alternatively, we note that the above ansatz is sufficient if $\alpha=r\pi$ for $r\in\mathbb{R}^{+}\backslash\mathbb{Q}^{+}$. Then $r$ can taken to be, for example, $r=q\pm\sqrt{2}\cdot 10^{-n}$, where $n$ can be arbitrarily large; Then we obtain a solution for the desired angle $\alpha=q\pi$ in the limit of large $n$.
IX Conclusions
In this paper, inspired by Moffatt eddies [25, 26] that emerge in wedge-like geometry under boundary forcing, we have shown that corner geometry with phoretically active sectors can be used to generate similar eddies. Our solution involved the exact solution of the diffusion problem in the wedge geometry using the Mellin transform formalism, followed by a solution of the biharmonic equation for the stream function of the associated flow. The proposed diffusiophoretic framework relies on a fixed-flux boundary condition for the solute release from a catalytic patch, which is appropriate when the diffusion of the solute is much faster than its advection (that is, in the limit of low Damköhler number).
In simple cases of uniform coverage of the walls with chemical activity, the geometric asymmetry alone is responsible for heterogeneous distribution of solute concentration, which drives the flow. In this case, the problem admits a simple analytical solution akin to the classical Taylor’s scraper problem when one wall is chemically active. Similarly, an analytical solution is also available when both walls are covered with catalyst. These solutions can serve as benchmarks for numerical computations, particularly when solving for the concentration field with non-trivial phoretic boundary conditions.
In the more general case, where patches of activity are distributed on the walls, we presented a solution technique using the Mellin transform to solve the diffusion problem and, subsequently, find the coefficients of the general solution of the associated flow problem. The solutions can be evaluated numerically for an arbitrary distribution of patches and combination of boundary conditions.
Funding This research was supported by the National Science Centre of Poland Sonata Bis grant no. 2023/50/E/ST3/00465 to ML.
Data availability statement The Mathematica notebooks used to compute the results of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.16938259, reference number 16938259.
Author ORCIDs M. Lisicki, https://orcid.org/0000-0002-6976-0281; D. Nowak, https://orcid.org/0009-0001-8334-2860